OpenStudy (anonymous):

Solving Systems Algebraically with Substitution x + 3y = 5 -2x - 4y = -5 I'm really confused on what I should do first? I think I solve for x which means minusing -3y to the other side of the equal sign... x = -3y + 5 Is that correct? :c

OpenStudy (anonymous):

yes that is right. but it wont be necessary to do it with the second equation just substitute the first equation that you get ( x = -3y + 5) to the x variable of the second equation, then just simplify it to solve for y. 2nd equation you have: -2x - 4y = -5 substitute x = -3y + 5 into the second equation: -2(-3y+5)-4y=-5 then solve for y

OpenStudy (anonymous):

Okay so it would go like this; 6y + 5 - 4y = -5 6y - 4y = -10 2y = -10 y = -5

OpenStudy (anonymous):

almost right! just a little mistake!